# Beyond Scale: The Diversity Coefficient as a Data Quality Metric for Variability in Natural Language Data

### v3

@@ Abstract
Keywords: data-centric machine learning, data quality, foundation models, pre-training, data diversity, diversity, large language models, LLMs
@@ 1 Introduction
Current trends in pre-training Large Language Models (LLMs) tend to concentrate on model and dataset size scaling Chowdhery et al. (2022); Nostalgebraist (2022); OpenAI (2023); Google (2023). Therefore, vast amounts of effort have been invested in understanding neural scaling laws—the power-law relationship between the loss of artificial deep networks and the size of the pre-training dataset for a fixed compute budget (Hestness et al., 2017; Rosenfeld et al., 2019; Henighan et al., 2020; Kaplan et al., 2020; Gordon et al., 2021; Hernandez et al., 2021; Jones, 2021; Zhai et al., 2022; Hoffmann et al., 2022; Clark et al., 2022; Neumann and Gros, 2022). In addition, recent work focuses on training a fixed-size model but using very large, trillion-token datasets (Touvron et al., 2023a, b). However, the effectiveness of these systems also fundamentally relies on the quality Longpre et al. (2023), variability and coverage of the pre-training data Hashimoto (2021); David et al. (2010) and not only the size. In particular, the richness and variety of data, otherwise known as data diversity, is a key aspect of data quality that plays an important role in researchers’ and practitioners’ choice of pre-training corpora for general capabilities (Gao et al., 2020; Brown et al., 2020; Touvron et al., 2023a; Eldan and Li, 2023; Gunasekar et al., 2023). In other words, diverse data is high quality data when your goal is to instill general capabilities in a model.
In addition, experimental settings demonstrate that the level of variety and diversity in pre-training data is likely a strong causal factor in the development of in-context learning (ICL) in LLMs, an essential component of the versatility and generality of LLMs (Xie et al., 2022; Shin et al., 2022; Chan et al., 2022). However, data quality, diversity and coverage David et al. (2010) are often discussed in vague and imprecise ways Longpre et al. (2023). Even in instances of quantitative approaches to data quality, methods are often not interpretable as to why certain datasets are preferable over others, or require reference corpora and, thereby, limit the diversity of resulting datasets (Xie et al., 2023a, c). Hence, we propose to ground the discussion of data quality and, in particular, data diversity, through the Task2Vec diversity coefficient for natural language—a metric using the expected distance between Task2Vec embeddings of data batches (Achille et al., 2019a; Miranda et al., 2022a) to quantify the level of structural and semantic diversity of natural language data, allowing researchers and practitioners to move beyond scale alone.
@@ 2.3 Backbone Used to compute the Diversity Coefficient
To compute Task2Vec embeddings, we use GPT-2 Radford et al. (2019) pre-trained on the English language as the probe network $f_{w}$. Following Task2Vec (Achille et al., 2019a), we fine-tune only the final layer (a language modeling head) on each batch since it’s the only tested method for computing Task2Vec embeddings Achille et al. (2019a); Miranda et al. (2022a, 2023), e.g. it’s not known if the intuitive properties observed in (Achille et al., 2019a) hold without fine-tuning the backbone. See Figure 1 for a visual of the diversity coefficient computation pipeline.
@@ 4 Related Work
A recently proposed diversity metric that doesn’t rely on an embedding function is the Vendi Score (Friedman and Dieng, 2022), which is the exponential of the Shannon entropy of the eigenvalues of a similarity matrix/kernel. For more discussion of related work, please see Appendix B.
@@ 5 Discussion
One potential limitation of our method is the need for a data representation. Although the requirement for a data representation might seem restrictive, we argue that it is an inherent aspect of data processing. Choosing symbols (e.g., one-hot vectors), raw pixels, etc. is a choice of data representation. We suggest deep learning representations due to their overwhelming success in machine learning, e.g. in computer vision (Krizhevsky et al., 2012; He et al., 2015), natural language processing (Devlin et al., 2018; Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Google, 2023), game playing (Silver et al., 2016; Mnih et al., 2013; Ye et al., 2021), theorem proving (Rabe et al., ; Polu and Sutskever, 2020; Han et al., ), code (Chen et al., ) and more. In addition, widely available open-source pre-trained models (e.g. CLIP (Radford et al., 2021), LLaMA (Touvron et al., 2023a), etc.) have made choosing a good embedding method easier. Another potential limitation of the diversity coefficient is the need to fine-tune a probe-network. Though this does introduce some computational overhead, it is relatively small (fine-tuning only the final layer) and outweighed by the utility of using information-rich Task2Vec embeddings.
Broader Impact Statement
This paper presents work whose goal is to advance the field of machine learning, and, in particular, presents tools for better understanding and training large language models (LLMs) via a rigorous approach to data quality. One important societal consequence of this is the fact that a better understanding of data quality would allow researchers and practitioners to make more informed decisions on dataset selection and more efficiently train LLMs. Thus, by reducing the need for the scaling of data quantity, this would likely lead to lower model development and inference costs, thereby making the development and use of LLMs more affordable and widely accessible to researchers and practitioners. Furthermore, more efficient training would lead to lower energy expenditure during model development, partially ameliorating any environmental degradation resulting from energy production and use. Finally, a rigorous understanding of data quality yields greater transparency about the value of data corpora, thereby allowing for fairer and better-grounded trade and arbitration with regard to the buying and selling of corpora. Aside from these factors, there are surely many other potential societal and ethical consequences of our work, but there are no others which we feel must be specifically highlighted here.
Acknowledgments and Disclosure of Funding
We want to acknowledge Rylan Schaeffer for discussions that helped us think more critically about our work. We would like to thank Simran Arora for her feedback on our manuscript. This research was funded in part by the Stanford School of Engineering fellowship and EDGE scholar fellowship from Stanford University. Alycia Lee is supported by Microsoft.
@@ Appendix A Intuition of Task2Vec
To better understand the Task2Vec embedding, observe that the (diagonal) of the FIM can be interpreted as a measure of the information that a given parameter contains about the generative distribution $p_{w}(\hat{x}_{t}\mid x_{t-1:1})$. Therefore, it serves as a unique fingerprint, or feature vector, for a batch, which defines a task distribution; as such, this vector, in its entirety, is as large as the number of parameters of the probe network. Empirical findings in (Achille et al., 2019a) show that Task2Vec embeddings cluster in a way that reflects semantics between different visual concepts, and that Task2Vec cosine distances are positively correlated with taxonomical distances.
@@
$E_{x\leavevmode\nobreak\ \sim\hat{p}}[KL(p_{w}^{\prime}||p_{w})]=\delta_{w}F\delta_{w}^{\top}+o(\delta(w)^{2})$
@@
Returning to the utility of a practical diversity metric, there is strong empirical support from efforts at increasing multi-modal image and text data quality that using semantically deduplicated data (as opposed to simply exact match deduplication) for training can yield impressive training efficiency gains and even improve out of domain generalization (Abbas et al., 2023, 2024). Though motivated by the same underlying hypothesis about the importance of diversity in data quality, our work on the diversity coefficient focuses on text-only data, the standard for training LLMs, and yields a more interpretable metric that allows practitioners to rigorously understand the quality of datasets, enabling the informed choice of baseline higher quality datasets rather than only being able to improve a chosen dataset. Furthermore, the generality of the diversity coefficient allows it to assess and inform the results of traditional data pruning, which has been shown to be a promising method for increasing data quality (Sorscher et al., 2023), while also lending itself naturally as a useful metric for guiding diverse data corpora construction, which has also been seen to be a fruitful endeavor with the release large, diverse, open source corpora (Gao et al., 2020).
@@
However, the benefits of this more sophisticated aggregation method are not clear, and its computation ($O(n^{3})$) is more expensive than the diversity coefficient ($O(n^{2})$). Moreover, it assumes a suitable similarity function/kernel, and does not provide guidance on data representation, arguably the most important ingredient in machine learning. Furthermore, they suggest that utilizing data representational methods such as embedding networks that require pretrained models may be limiting. We argue instead that data representation is a fundamental property of data processing that has led to the overwhelming success in machine learning due to deep learning, e.g. in computer vision (Krizhevsky et al., 2012; He et al., 2015), natural language processing (Devlin et al., 2018; Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Google, 2023), game playing (Silver et al., 2016; Mnih et al., 2013; Ye et al., 2021), theorem proving (Rabe et al., ; Polu and Sutskever, 2020; Han et al., ), code (Chen et al., ) and more. For further comments on the Vendi Score see section B.
@@ C.2 Token Based Definitions of Diversity are Brittle to Paraphrasings
One can also consider formalizations of diversity based on token distributions in a batch (or data set). We did not use this because previous work has already demonstrated its limitation as data representations for train-test data contamination detection (Jiang et al., 2024). This is highly related to our work because contamination detection requires good data representations for detecting differences in text. In particular, they say: “these contamination definitions primarily revolve around n-gram or token overlaps, which only target direct duplications present in both training and evaluation datasets and might provide both high false positive rate (since many semantically different texts have overlaps) and false negative rate (since simple paraphrasing can evade detection)”. To make it more concrete, consider an example. Consider the sentences “The great paper” and “the fantastic paper”. These would be binned to entirely different locations in a histogram, while the difference in the sentence is varying much more smoothly. Therefore, token based representation of data are brittle to simple paraphrasing and as our example shows result in overly “harsh” metrics—similar to previous work that criticizes harsh metrics for LLM evaluation (Schaeffer et al., 2023a). In addition, previous work has shown that importance weights of data like text are statistically intractable to estimate without sufficient additional structure (Xie et al., 2023b; Bengtsson et al., 2008; Gelman and Meng, 2004; Snyder et al., 2008). For example, the histograms of different types of text end up being sparse, even with hashing e.g., consider 50k (GPT2), 30K (LLaMA2) or 10K (DSIR) feature vectors. We conjecture that if these representations were robust, the transformer architecture (Vaswani et al., 2017) wouldn’t have emerged as the successful architecture that it is today (2024).
@@ C.3 Discussion on Model Agnostics Formalizations of Data Diversity
Theoretical Relation of FIM (Model) and Kolmogorov (Model Agnostic) based metric: Theory can be important to have a better understanding of the world. Therefore, theoreticians might be interested in diversity definitions that are model agnostic, since, they might be interested in the inherit diversity of the data or task. We emphasize that Task2Vec was derived with such considerations in mind: “The FIM is also related to the (Kolmogorov) complexity of a task, a property that can be used to define a computable metric of the learning distance between tasks” (Achille et al., 2019a). More precisely (Achille et al., 2018) showed that the training dynamics of a deep network minimizing a loss function can be viewed as approximately optimizing an upper bound on the Kolmogorov Structure Function (KSF) of the learning task. This connection arises because the loss function incorporates a complexity term measured by the Fisher Information Matrix of the network weights. Therefore, Fisher information provides a computable way to relate the geometry of the loss landscape to notions of task complexity linked to Kolmogorov complexity. Therefore, Task2Vec (via FIM) was derived with Kolmogorov Complexity in mind. In Summary, the relationship between Fisher Information Matrix (FIM) and Kolmogorov Complexity, as discussed in Achille’s works (Achille et al., 2019a, 2018, b, 2021), suggests that FIM can serve as an upper bound for Kolmogorov Complexity in the context of learning tasks. Specifically, the FIM captures the sensitivity of the learning model to its parameters, reflecting the amount of information about the data that the model encodes. This sensitivity can indirectly bound the Kolmogorov Complexity by quantifying the model’s complexity and its capacity to learn from data, thus offering a bridge between empirical measures of complexity (FIM) and theoretical ones (Kolmogorov Complexity). However, explicit relationships quantifying how distances based on Fisher information matrices correlate with true Kolmogorov complexity are not provided.
@@ G.2 Higher Formal Diversity of the Training Data Set Causes Higher Test Performance: All Results
Figure 7: Despite smaller scale and relatively high loss/perplexity on certain evaluation datasets, the experimental models still grasp important aspects of sentence syntax and structure, even on evaluation data they otherwise struggle on (e.g. Wikitext-103, for which all models scored around 6.6 loss or more). In addition, when evaluated on similar, well aligned datasets (e.g. USPTO + PubMed Abs. (train) on USPTO (validation)), models are sometimes also able to match the form and semantic content of the example. Input examples were randomly picked from the given evaluation dataset.
@@ I.6 Details on Mixes used for Interleaved Datasets
Importantly, the relationship between pre-training dataset diversity and test performance remains a key question. To investigate, we conduct preliminary experiments, pre-training three GPT-2 models on datasets of varying diversities, then evaluate their performance on diverse validation datasets. We observe in Table 2 that increased diversity leads to decreased cross-entropy loss, indicating a positive relationship between diversity and model performance on diverse tasks. Although the cross-entropy values are arguably large, Figure 7 shows that models were still pre-trained enough to grasp important aspects of sentence syntax and structure of their respective dataset, indicating model performance on evaluation represents meaningful improvements in ability. Hence, we conjecture pre-training on higher diversity data improves test performance on diverse tasks, though more extensive experiments are needed to know so conclusively.

### v4

@@ 1 Introduction
Current trends in pre-training Large Language Models (LLMs) tend to concentrate on model and dataset size scaling Chowdhery et al. (2022); Nostalgebraist (2022); OpenAI (2023); Google (2023). Therefore, vast amounts of effort have been invested in understanding neural scaling laws—the power-law relationship between the loss of artificial deep networks and the size of the pre-training dataset for a fixed compute budget (Hestness et al., 2017; Rosenfeld et al., 2019; Henighan et al., 2020; Kaplan et al., 2020; Gordon et al., 2021; Hernandez et al., 2021; Jones, 2021; Zhai et al., 2022; Hoffmann et al., 2022; Clark et al., 2022; Neumann & Gros, 2022). In addition, recent work focuses on training a fixed-size model but using very large, trillion-token datasets (Touvron et al., 2023a; b). However, the effectiveness of these systems also fundamentally relies on the quality Longpre et al. (2023), variability and coverage of the pre-training data Hashimoto (2021); David et al. (2010) and not only the size. In particular, the richness and variety of data, otherwise known as data diversity, is a key aspect of data quality that plays an important role in researchers’ and practitioners’ choice of pre-training corpora for general capabilities (Gao et al., 2020; Brown et al., 2020; Touvron et al., 2023a; Eldan & Li, 2023; Gunasekar et al., 2023). In other words, diverse data is high quality data when your goal is to instill general capabilities in a model.
In addition, experimental settings demonstrate that the level of variety and diversity in pre-training data is likely a strong causal factor in the development of in-context learning (ICL) in LLMs, an essential component of the versatility and generality of LLMs (Xie et al., 2022; Shin et al., 2022; Chan et al., 2022). However, data quality, diversity and coverage David et al. (2010) are often discussed in vague and imprecise ways Longpre et al. (2023). Even in instances of quantitative approaches to data quality, methods are often not interpretable as to why certain datasets are preferable over others, or require reference corpora and, thereby, limit the diversity of resulting datasets (Xie et al., 2023a; c). Hence, we propose to ground the discussion of data quality and, in particular, data diversity, through the Task2Vec diversity coefficient for natural language—a metric using the expected distance between Task2Vec embeddings of data batches (Achille et al., 2019a; Miranda et al., 2022a) to quantify the level of structural and semantic diversity of natural language data, allowing researchers and practitioners to move beyond scale alone.
@@ 2.3 Backbone Used to compute the Diversity Coefficient
To compute Task2Vec embeddings, we use GPT-2 Radford et al. (2019) pre-trained on the English language as the probe network $f_{w}$. Following Task2Vec (Achille et al., 2019a), we fine-tune only the final layer (a language modeling head) on each batch since it’s the only tested method for computing Task2Vec embeddings Achille et al. (2019a); Miranda et al. (2022a; 2023), e.g. it’s not known if the intuitive properties observed in (Achille et al., 2019a) hold without fine-tuning the backbone. See Figure 1 for a visual of the diversity coefficient computation pipeline.
@@ 4 Related Work
A recently proposed diversity metric that doesn’t rely on an embedding function is the Vendi Score (Friedman & Dieng, 2022), which is the exponential of the Shannon entropy of the eigenvalues of a similarity matrix/kernel. For more discussion of related work, please see Appendix B.
@@ 5 Discussion
One potential limitation of our method is the need for a data representation. Although the requirement for a data representation might seem restrictive, we argue that it is an inherent aspect of data processing. Choosing symbols (e.g., one-hot vectors), raw pixels, etc. is a choice of data representation. We suggest deep learning representations due to their overwhelming success in machine learning, e.g. in computer vision (Krizhevsky et al., 2012; He et al., 2015), natural language processing (Devlin et al., 2018; Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Google, 2023), game playing (Silver et al., 2016; Mnih et al., 2013; Ye et al., 2021), theorem proving (Rabe et al., ; Polu & Sutskever, 2020; Han et al., ), code (Chen et al., ) and more. In addition, widely available open-source pre-trained models (e.g. CLIP (Radford et al., 2021), LLaMA (Touvron et al., 2023a), etc.) have made choosing a good embedding method easier. Another potential limitation of the diversity coefficient is the need to fine-tune a probe-network. Though this does introduce some computational overhead, it is relatively small (fine-tuning only the final layer) and outweighed by the utility of using information-rich Task2Vec embeddings.
@@ Appendix A Intuition of Task2Vec
To better understand Task2Vec embeddings, observe that the (diagonal) of the FIM can be interpreted as a measure of the information that a given parameter contains about the generative distribution $p_{w}(\hat{x}_{t}\mid x_{t-1:1})$. Therefore, it serves as a unique fingerprint, or feature vector, for a batch, which defines a task distribution; as such, this vector, in its entirety, is as large as the number of parameters of the probe network. Empirical findings in (Achille et al., 2019a) show that Task2Vec embeddings cluster in a way that reflects semantics between different visual concepts, and that Task2Vec cosine distances are positively correlated with taxonomical distances.
@@
$E_{x~{}\sim\hat{p}}[KL(p_{w}^{\prime}||p_{w})]=\delta_{w}F\delta_{w}^{\top}+o(\delta(w)^{2})$
@@
Returning to the utility of a practical diversity metric, there is strong empirical support from efforts at increasing multi-modal image and text data quality that using semantically deduplicated data (as opposed to simply exact match deduplication) for training can yield impressive training efficiency gains and even improve out of domain generalization (Abbas et al., 2023; 2024). Though motivated by the same underlying hypothesis about the importance of diversity in data quality, our work on the diversity coefficient focuses on text-only data, the standard for training LLMs, and yields a more interpretable metric that allows practitioners to rigorously understand the quality of datasets, enabling the informed choice of baseline higher quality datasets rather than only being able to improve a chosen dataset. Furthermore, the generality of the diversity coefficient allows it to assess and inform the results of traditional data pruning, which has been shown to be a promising method for increasing data quality (Sorscher et al., 2023), while also lending itself naturally as a useful metric for guiding diverse data corpora construction, which has also been seen to be a fruitful endeavor with the release large, diverse, open source corpora (Gao et al., 2020).
@@
However, the benefits of this more sophisticated aggregation method are not clear, and its computation ($O(n^{3})$) is more expensive than the diversity coefficient ($O(n^{2})$). Moreover, it assumes a suitable similarity function/kernel, and does not provide guidance on data representation, arguably the most important ingredient in machine learning. Furthermore, they suggest that utilizing data representational methods such as embedding networks that require pretrained models may be limiting. We argue instead that data representation is a fundamental property of data processing that has led to the overwhelming success in machine learning due to deep learning, e.g. in computer vision (Krizhevsky et al., 2012; He et al., 2015), natural language processing (Devlin et al., 2018; Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Google, 2023), game playing (Silver et al., 2016; Mnih et al., 2013; Ye et al., 2021), theorem proving (Rabe et al., ; Polu & Sutskever, 2020; Han et al., ), code (Chen et al., ) and more. For further comments on the Vendi Score see section B.
@@ C.2 Token Based Definitions of Diversity are Brittle to Paraphrasings
One can also consider formalizations of diversity based on token distributions in a batch (or data set). We did not use this because previous work has already demonstrated its limitation as data representations for train-test data contamination detection (Jiang et al., 2024). This is highly related to our work because contamination detection requires good data representations for detecting differences in text. In particular, they say: “these contamination definitions primarily revolve around n-gram or token overlaps, which only target direct duplications present in both training and evaluation datasets and might provide both high false positive rate (since many semantically different texts have overlaps) and false negative rate (since simple paraphrasing can evade detection)”. To make it more concrete, consider an example. Consider the sentences “The great paper” and “the fantastic paper”. These would be binned to entirely different locations in a histogram, while the difference in the sentence is varying much more smoothly. Therefore, token based representation of data are brittle to simple paraphrasing and as our example shows result in overly “harsh” metrics—similar to previous work that criticizes harsh metrics for LLM evaluation (Schaeffer et al., 2023a). In addition, previous work has shown that importance weights of data like text are statistically intractable to estimate without sufficient additional structure (Xie et al., 2023b; Bengtsson et al., 2008; Gelman & Meng, 2004; Snyder et al., 2008). For example, the histograms of different types of text end up being sparse, even with hashing e.g., consider 50k (GPT2), 30K (LLaMA2) or 10K (DSIR) feature vectors. We conjecture that if these representations were robust, the transformer architecture (Vaswani et al., 2017) wouldn’t have emerged as the successful architecture that it is today (2024).
@@ C.3 Discussion on Model Agnostics Formalizations of Data Diversity
Theoretical Relation of FIM (Model) and Kolmogorov (Model Agnostic) based metric: Theory can be important to have a better understanding of the world. Therefore, theoreticians might be interested in diversity definitions that are model agnostic, since, they might be interested in the inherit diversity of the data or task. We emphasize that Task2Vec was derived with such considerations in mind: “The FIM is also related to the (Kolmogorov) complexity of a task, a property that can be used to define a computable metric of the learning distance between tasks” (Achille et al., 2019a). More precisely (Achille et al., 2018) showed that the training dynamics of a deep network minimizing a loss function can be viewed as approximately optimizing an upper bound on the Kolmogorov Structure Function (KSF) of the learning task. This connection arises because the loss function incorporates a complexity term measured by the Fisher Information Matrix of the network weights. Therefore, Fisher information provides a computable way to relate the geometry of the loss landscape to notions of task complexity linked to Kolmogorov complexity. Therefore, Task2Vec (via FIM) was derived with Kolmogorov Complexity in mind. In Summary, the relationship between Fisher Information Matrix (FIM) and Kolmogorov Complexity, as discussed in Achille’s works (Achille et al., 2019a; 2018; b; 2021), suggests that FIM can serve as an upper bound for Kolmogorov Complexity in the context of learning tasks. Specifically, the FIM captures the sensitivity of the learning model to its parameters, reflecting the amount of information about the data that the model encodes. This sensitivity can indirectly bound the Kolmogorov Complexity by quantifying the model’s complexity and its capacity to learn from data, thus offering a bridge between empirical measures of complexity (FIM) and theoretical ones (Kolmogorov Complexity). However, explicit relationships quantifying how distances based on Fisher information matrices correlate with true Kolmogorov complexity are not provided.
@@ G.2 Higher Formal Diversity of the Training Data Set Causes Higher Test Performance: All Results
Figure 7: Despite smaller scale (51M parameters) and relatively high loss/perplexity on certain evaluation datasets, the experimental models still grasp important aspects of sentence syntax and structure, even on evaluation data they otherwise struggle on (e.g. Wikitext-103, for which all models scored around 6.6 loss or more). In addition, when evaluated on similar, well aligned datasets (e.g. USPTO + PubMed Abs. (train) on USPTO (validation)), models are sometimes also able to match the form and semantic content of the example. Input examples were randomly picked from the given evaluation dataset.
@@ I.6 Details on Mixes used for Interleaved Datasets
Importantly, the relationship between pre-training dataset diversity and test performance remains a key question. To investigate, we conduct preliminary experiments, pre-training three GPT-2 models on datasets of varying diversities, then evaluate their performance on diverse validation datasets. We observe in Table 2 that increased diversity leads to decreased cross-entropy loss, indicating a positive relationship between diversity and model performance on diverse tasks. Although the cross-entropy values are arguably large, Figure 7 shows that even the smallest 51M parameter models were still pre-trained enough to grasp important aspects of sentence syntax and structure of their respective dataset, indicating model performance on evaluation represents meaningful improvements in ability. Hence, we conjecture pre-training on higher diversity data improves test performance on diverse tasks, though more extensive experiments are needed to know so conclusively.